# When Judgment Becomes Noise: How Design Failures in LLM Judge Benchmarks Silently Undermine Validity

### v2

@@ 2 Methods
Our chosen measure is a simple aggregate of internal consistency and discriminant validity: Cronbach’s $\alpha$ for the former, and a cross-loading ratio (CLR) with sigmoid normalization centered at 1.5, and HTMT for the latter, formally –
$R_{\text{psychometric}}$ $=\tfrac{1}{3}\Big(\tfrac{1}{k}\sum_{j=1}^{k}\alpha_{j}\;+\tfrac{1}{k}\sum_{j=1}^{k}\text{CLR}_{\text{norm},j}\;+\big(1-\tfrac{2}{k(k-1)}\sum_{i<j}\text{HTMT}_{ij}\big)\Big),$ (4)
with sensitivity $\text{Sens}_{\text{psychometric}}=\sqrt{1-R_{\text{psychometric}}}\times\text{score\_range}$. Full definitions are provided in Section B.2.
@@ 4 Results
Figure 4: ELO-style aggregation compresses multi-dimensional, noisy judgments into apparently smooth rankings, masking upstream uncertainty.
@@ B.2 Psychometric Validity
  • Heterotrait-Monotrait Ratio (HTMT): Assesses construct validity between factors
@@
The HTMT ratio between factors $f_{i}$ and $f_{j}$ is: $\text{HTMT}_{ij}=\frac{\bar{r}_{ij}}{\sqrt{\bar{r}_{ii}\cdot\bar{r}_{jj}}}$ where $\bar{r}_{ij}$ is the mean correlation between items of different factors, and $\bar{r}_{ii}$ is the mean correlation between items within factor $i$.
Definition 9 (Sigmoid CLR Normalization).
To address asymmetric penalties in CLR assessment while maintaining literature-supported thresholds Campbell & Fiske (1959), we apply sigmoid normalization centered at the psychometric threshold CLR = 1.5 from the 0.40-0.30-0.20 rule Hair et al. (2019), which states that items should have strong factor loadings on their intended factors and relatively low loadings on all secondary factors: $\text{CLR}_{\text{norm},i}=\frac{1}{1+\exp(-2(\text{CLR}_{i}-1.5))}$ This transformation provides smooth, bounded scoring where CLR = 1.5 yields 0.5, poor discriminant validity (CLR $\approx$ 0) yields $\approx$ 0.05, and strong validity (CLR $\geq$ 2.0) yields $\geq$ 0.73.
@@
We combine these metrics into a unified psychometric validity score using equal weighting:
$R_{\text{psychometric}}=\frac{1}{3}\left(\frac{1}{k}\sum_{i=1}^{k}\alpha_{i}+\frac{1}{k}\sum_{i=1}^{k}\text{CLR}_{\text{norm},i}+\left(1-\frac{2}{k(k-1)}\sum_{i<j}\text{HTMT}_{ij}\right)\right)$
where the HTMT component is inverted since lower HTMT values indicate better discriminant validity.
@@
The sigmoid CLR normalization addresses several methodological concerns identified in our analysis:
@@
  • Balanced weighting: Prevents poor CLR values from creating excessive negative penalties that dominate well-established metrics (Cronbach’s $\alpha$, HTMT)

### v3

@@ 2 Methods
We aggregate internal consistency and discriminant validity using normalized, thresholded components and an explicit penalty for judge failure rates. Let $\alpha_{i}$ denote Cronbach’s alpha for factor $i$, $\text{CLR}_{i}$ its cross-loading ratio, $\overline{\text{HTMT}}_{i\cdot}$ the mean HTMT between factor $i$ and all others (computed on absolute item correlations), and $\phi_{i}\in[0,1]$ the fraction of missing/unscorable judgments for factor $i$. Define
$\alpha_{i}^{\mathrm{norm}}=\tfrac{\alpha_{i}-0.70}{0.95-0.70}$
$\text{CLR}_{i}^{\mathrm{norm}}=\tfrac{\text{CLR}_{i}-1.0}{2.0-1.0}$
$\text{HTMT}i^{\mathrm{norm}}=1-\tfrac{1}{0.85}$
Let $D_{i}$ be the harmonic mean of the two discriminant components, $D_{i}=\tfrac{2\,\text{CLR}^{\mathrm{norm}}_{i}\,\text{HTMT}^{\mathrm{norm}}_{i}}{\text{CLR}^{\mathrm{norm}}_{i}+\text{HTMT}^{\mathrm{norm}}_{i}}$. The per-factor validity is $V_{i}=\tfrac{1}{2}\alpha^{\mathrm{norm}}_{i}+\tfrac{1}{2}D_{i}$ and we apply a failure penalty multiplicatively, $\tilde{V}_{i}=(1-\phi_{i})V_{i}$. Our overall psychometric validity is the mean across factors,
$R_{\text{psychometric}}=\tfrac{1}{k}\sum_{i=1}^{k}\tilde{V}_{i},$ (4)
with sensitivity $\text{Sens}_{\text{psychometric}}=\sqrt{1-R_{\text{psychometric}}}\times\text{score\_range}$. Full definitions appear in Section B.2.
@@ 4 Results
Psychometric validity is effective for distinguishing quality judgment components from problematic ones.
Using the validity formulation in Section B.2, we reevaluated Setting 1 across four judges (DeepSeek-R1-32B, GPT-3.5-Turbo-0125, GPT-4o-mini-0718, QwQ-32B). Figure 2 summarizes internal consistency (Cronbach’s $\alpha$), discriminant validity (HTMT computed on absolute correlations; CLR normalized to $[0,1]$ around a 1.0/2.0 baseline/target), and the resulting overall validity score (harmonic mean of discriminant subcomponents blended with $\alpha$, penalized by the factor-wise failure rate).
Completeness, correctness, and style attain excellent internal consistency across judges ($\alpha\gtrsim 0.90$) and moderate discriminant signal. This suggests both that the Arena-Hard-Auto benchmark may be suitable for judging such factors, but also that judges frequently confuse them.
Safety is systematically weaker: internal consistency is lower for some judges and the factor shows elevated failure rates (fraction of missing/unscorable judgments), which directly reduces validity. The most extreme outlier is GPT-3.5 on safety (largely due to failures of the judge to respond); this drives down the validity score despite otherwise reasonable separation on the remaining factors.
Conciseness shows high validity – unfortunately, as we know from factor analysis, this is in part because it is ignored in the overall judgment.
Figure 2: Psychometric validity summary for Setting 1 across four judges. Bars show (top-left) Cronbach’s $\alpha$ (internal consistency); (top-right) cross-loading ratio (CLR; higher indicates stronger factor separation, shown also as a normalized $[0,1]$ score in the validity computation); (bottom-left) HTMT computed on absolute item-level correlations (lower is better; the dashed line marks the 0.85 threshold); and (bottom-right) overall validity score combining $\alpha$ and discriminant components via a harmonic mean and applying a multiplicative penalty $(1-\phi)$ for factor-wise failure rate $\phi$ (share of unscorable judgments such as “Safety: N/A”). Atypical results appear primarily on the safety factor for GPT-3.5, where high failure rates depress the overall score.
@@
Figure 4: ELO-style aggregation compresses multi-dimensional, noisy judgments into apparently smooth rankings, masking upstream uncertainty. Even more so than in general-case factor analysis, the feature loadings under ELO exhibit clear and strong positive or negative correlation.
@@ B.2 Psychometric Validity
  • Heterotrait-Monotrait Ratio (HTMT): Assesses construct validity between factors (computed on absolute item–item correlations)
@@
The HTMT ratio between factors $f_{i}$ and $f_{j}$ is:
$\text{HTMT}_{ij}=\frac{\overline{|r|}_{ij}}{\sqrt{\overline{|r|}_{ii}\cdot\overline{|r|}_{jj}}}$
where $\overline{|r|}_{ij}$ is the mean absolute correlation between items of different factors, and $\overline{|r|}_{ii}$ is the mean absolute correlation between items within factor $i$. Using absolute correlations matches the conventional HTMT definition and avoids sign artefacts.
Definition 9 (Bounded CLR Normalization).
We map the cross-loading ratio to $[0,1]$ with a simple bounded linear transform using 1.0 as the “no separation” baseline and 2.0 as a strong separation target supported by practice Hair et al. (2019):
$\text{CLR}_{\text{norm},i}=\mathrm{clip}\big((\text{CLR}_{i}-1.0)/(2.0-1.0),\,0,\,1\big).$
This avoids asymmetric penalties and makes the scale interpretable: CLR $\leq$ 1.0 $\Rightarrow$ 0, CLR $\geq$ 2.0 $\Rightarrow$ 1.
Definition 10 (Failure Rate).
For each factor $f_{i}$, define a failure rate $\phi_{i}\in[0,1]$ as the fraction of judgments that are missing or unscorable for that factor (e.g., rubric omissions such as “Safety: N/A”). This term down-weights validity when a factor is rarely applied by a judge.
@@
Let $\mathrm{clamp}(x;a,b)=\min\{\max\{x,a\},b\}$ and $H(a,b)=\tfrac{2ab}{a+b}$ denote the harmonic mean. Define per-factor components
$\alpha^{\mathrm{norm}}_{i}$ $=\mathrm{clamp}\Big(\frac{\alpha_{i}-0.70}{0.95-0.70},\,0,\,1\Big),$
$\text{CLR}^{\mathrm{norm}}_{i}$ $=\mathrm{clamp}\Big(\frac{\text{CLR}_{i}-1.0}{2.0-1.0},\,0,\,1\Big),$
$\text{HTMT}^{\mathrm{norm}}_{i}$ $=\mathrm{clamp}\Big(1-\frac{1}{0.85}\cdot\overline{\text{HTMT}}_{i\cdot},\,0,\,1\Big),$
where $\overline{\text{HTMT}}_{i\cdot}$ is the average HTMT between factor $i$ and all others. The discriminant component is
$D_{i}=H\big(\text{CLR}^{\mathrm{norm}}_{i},\,\text{HTMT}^{\mathrm{norm}}_{i}\big).$
The per-factor validity is then
$V_{i}=\tfrac{1}{2}\,\alpha^{\mathrm{norm}}_{i}+\tfrac{1}{2}\,D_{i},$
with a failure penalty applied multiplicatively: $\tilde{V}_{i}=(1-\phi_{i})\,V_{i}$. The overall psychometric validity aggregates across factors:
$R_{\text{psychometric}}=\frac{1}{k}\sum_{i=1}^{k}\tilde{V}_{i}.$
@@
The bounded CLR normalization and the redundancy-aware discriminant component address several methodological concerns:
@@
  • Balanced weighting: Uses the harmonic mean of CLR and HTMT normalizations so that disagreement between the two signals is penalized rather than double-counted
@@
B.3 Reliability Heatmaps, Failure Rates, and Distributions
Figure 6: Internal consistency (Cronbach’s $\alpha$) by judge and factor (Setting 1). Higher is better; values near 1 indicate highly consistent item-level judgments within a factor. Most factors attain $\alpha\geq 0.9$ across judges; safety is comparatively weaker for some judges.
Figure 7: Construct validity via HTMT (computed on absolute item correlations) by judge and factor. Lower is better; a conventional 0.85 threshold marks acceptable discriminant validity. Values near or above the threshold indicate factor overlap.
Figure 8: Failure rates $\phi$ (fraction of missing/unscorable judgments) by judge and factor. This term down-weights validity: $\tilde{V}=(1-\phi)V$. The highest rates occur for GPT-3.5 on safety, reflecting frequent rubric omissions (e.g., “Safety: N/A”).
Figure 9: Distribution of reliability/validity components across factors. Violin plots summarize dispersion of Cronbach’s $\alpha$, CLR, HTMT, and the composite validity score. Narrow, high-location violins indicate strong, stable performance; wide, tall, or low-location violins flag potential instability.